# Flow with FlorDB: Incremental Context Maintenance for the Machine Learning Lifecycle

### v1

@@ 2 N = get_num_pages(doc_name)
3.2. Managing Dataflow and Feedback Loops
As we have discussed throughout this paper, context is something you log, and it’s best when it can be logged post-hoc and on demand. But also, and more importantly, context is something you build and maintain. “Build Context” is information about how data is created and used: e.g., dependency management for distribution and building across different machines and by different people; provenance and lineage; routes, pathways, or branches in pipelines; and flow of control and data. “Change Context” is information about the version history of data, code, configuration parameters, checkpoints, and associated information, including changes over time to both structure and content Hellerstein et al. (2017). Build and Change Context is what allows us to answer important development questions such as the following: Where did you define this data? Where do you transform it? Who was the last person to change this? Where should I make changes to add another transformation or filter? And so on.
Currently, these questions are often answered through face-to-face interactions or via communication tools like Slack or email Shankar et al. (2024). However, this approach has several limitations, primarily: i) friction, inefficiency and possible delays, as it relies on the availability and memory of colleagues, and ii) lack of scalability, especially as teams change or projects evolve. Alternatively, developers may try to anticipate frequently asked questions (FAQs) and maintain comprehensive READMEs, documentation, wikis, or forums. However, as previously noted, this level of discipline and methodology can conflict with the high rate of iteration and agility typical in MLOps environments.
FlorDB addresses these challenges by managing context effectively. When a colleague has questions about the project, they can simply invoke a flor.dataframe with the desired data (originating from flor.log(name, value)). This dataframe automatically gathers the requisite context alongside the queried data. In essence, FlorDB is a step towards making work self-documenting, a direct consequence of effective context management. By logging with FlorDB, the MLE can build an ML pipeline, and by building that pipeline, a user or colleague can navigate versions, answer questions about provenance or dataflow, identify the files and lines where data is entered and transformed, and so on.
Figure 3 illustrates an ML pipeline for a document intelligence application with feedback loops defined in a Makefile. The pipeline consists of three main components:
  1. (1)
    infer: Uses flor.dataframe("acc", "recall") to select the model checkpoint with the highest recall (or a fallback model if no checkpoint exists). The corresponding dataframe queried by infer is shown on the right.
  2. (2)
    run: Launches a Flask web application to display model predictions, overlaying a color layer (denoting polarity) over each page for human review. Upon confirmation, Flask logs the final color of each page, incorporating feedback into the cycle.
  3. (3)
    train: Utilizes the human-reviewed data (feedback) to fine-tune the model and improve its performance for future rounds of document analysis (Figure 4).
This process is iterative, alternating between calls to make run and make train. Throughout this cycle, flor.log(name, value) manages writes, while flor.dataframe(*args) handles reads, ensuring efficient data flow and context management.
1process_pdfs: \$(PDFS) pdf_demux.py
2 @echo "Processing PDF files..."
3 @python pdf_demux.py
4 @touch process_pdfs
5
6featurize: process_pdfs featurize.py
7 @echo "Featurizing Data..."
8 @python featurize.py
9 @touch featurize
10
11train: featurize hand_label train.py
12 @echo "Training..."
13 @python train.py
14
15model.pth: train export_ckpt.py
16 @echo "Generating model..."
17 @python export_ckpt.py
18
19infer: model.pth infer.py
20 @echo "Inferencing..."
21 @python infer.py
22 @touch infer
23
24hand_label: label_by_hand.py
25 @echo "Labeling by hand"
26 @python label_by_hand.py
27 @touch hand_label
28
29run: featurize infer
30 @echo "Starting Flask..."
31 @flask run
4. PDF Parser Demo
This section gives an overview of the PDF Parser Demo, a practical application of FlorDB in document intelligence. We use FlorDB to manage context and dataflow, and demonstrate how it can be effectively used to implement and manage a real-world AI/ML applications with ML pipelines of moderate complexity. The PDF Parser [^2] https://github.com/ucbepic/pdf_parser is a Flask-based web application designed for efficient PDF document processing, including tasks like splitting PDFs, extracting text, and preparing data for analysis using natural language processing (NLP) techniques. Users interact with the parser through a simple web interface, as shown in Figure 5 (left). This interface allows users to navigate PDF documents and select the desired processing options. Beyond demonstrating the practical value of a context-rich approach, we hope it will serve as a reference implementation for those looking to get started with FlorDB.
4.1. PDF Extraction & Text Featurization
Once the PDF is converted into text and image formats, ensuring there is one document per page, the process of featurization begins (Figure 2). This process typically involves text extraction, feature engineering, and vectorization. This featurization process is essential for transforming raw PDF data into a structured form that is amenable to analysis and machine learning applications. The described methodology focuses on maximizing the information extracted from each page, ensuring that both textual and visual data contribute to the inferences made on the document. Takeaway: When used in featurization contexts (Figure 2), FlorDB can morph into a feature store.
4.2. Inference Pipeline
The inference pipeline automates the processing and analysis of images organized into document-specific folders. The inference pipeline queries FlorDB to select a model checkpoint, and uses it to systematically processes each document’s pages and images. For each image, it loads the image file, applies standard pre-processing steps such as resizing and normalization, and converts it into a format suitable for input to the pre-trained model. The pre-trained model is chosen based on user-defined validation metrics. The model then makes predictions on the images, logging the final inferences. Takeaway: When used in inference pipelines, FlorDB can morph into a model registry (dataframe in Figure 3) for checkpoint selection.
4.3. Training Pipeline
The training pipeline encapsulates a typical machine learning workflow, tailored for classifying images extracted from PDF pages (Figure 4). This pipeline performs a load of training data (line 1 in Figure 4, managed by FlorDB), and data preparation. A model architecture is defined, and the model is trained over a number of epochs. Performance is monitored by flor logging, and context is managed and tracked using FlorDB. Takeaway: When used in training pipelines, FlorDB can morph into a training data store (line 1 in Figure 4), and a model repository (lines 3-21 in Figure 4).
@@ 5. Discussion
We assess FlorDB through the lens of the 3Vs of MLOps Shankar et al. (2024):
  • Velocity: FlorDB enhances the speed of ML model development, testing, and deployment by streamlining the integration of new data and models, enabling rapid iteration and adaptation.
@@
  • Versioning: FlorDB leverages version control for managing data, models, and code changes, supporting detailed tracking and allowing teams to understand model evolution over time.
@@ 6. Related Work
Experiment Tracking and Version Control: Systems like MLFlow Zaharia et al. (2018), DVC dvc (2021), and Weights & Biases focus on managing experiments and ensuring reproducibility. They provide tools for tracking experiments, packaging code, and sharing models. While crucial for tracking model and data evolution, they primarily concentrate on experiment management without deeply integrating hindsight logging capabilities.
@@
End-to-End ML Workflows: AWS SageMaker and Kubeflow provide comprehensive solutions for building, training, and deploying ML models Liberty et al. (2020); Community (2021). They offer tools for data labeling, model training, and hosting, supporting scalable ML workflows. Both platforms emphasize scalability and operational efficiency but primarily focus on deployment and operational aspects. FlorDB can use either system as a drop-in replacement to Make, the default build system. Other systems in this space include Helix Xin et al. (2018) and Motion Shankar and Parameswaran (2024).
@@ 7. Conclusion
FlorDB represents a step toward managing the complex metadata and context associated with the machine learning lifecycle. By addressing the enduring challenges of metadata management in AI/ML development, FlorDB offers a flexible yet powerful solution aligning with modern MLOps agile development practices.
The key contributions of FlorDB include several features that enhance the developer experience and streamline machine learning workflows. First, Incremental Context Maintenance allows developers to gradually build out the metadata and project structure within their existing workflows. Second, the extensions presented to support pipelines, dataflow, and feedback are backward compatible with Multiversion Hindsight Logging allowing for a “metadata later” approach, which enables the addition and refinement of metadata post-hoc. This capability supports rapid iteration and adaptation without compromising the long-term maintainability of projects. Last, FlorDB offers a Unified Framework & Versatility by integrating features from bespoke machine learning metadata systems. It provides a cohesive solution for managing the entire ML lifecycle, from data preparation to model deployment and feedback integration. For instance, as demonstrated through the PDF Parser application, FlorDB can take on multiple roles within the ML pipeline, such as acting as a feature store, model registry, training data store, and experiment record.

### v2

@@ 4.3. Training Pipeline
4.4. Closing the Loop: Feedback via UI
The main Flask script outlines the core functionalities of a web application designed for handling PDF documents and associated image files. It includes routes for displaying and manipulating PDFs and their converted image previews within a web interface, while also incorporating human-in-the-loop feedback for improving model performance. The core of the application is structured around Flask routes that handle web requests and facilitate expert feedback. Through the UI, domain experts can review model predictions and provide corrective labels, which are managed with the same metadata infrastructure as computational steps. This human feedback loop is crucial for iteratively improving model performance and maintaining data quality.
Helper functions such as get_colors() fetch color data associated with the pages of a document, integrating both automated predictions and human-provided labels from the dataset (line 6 in Figure 6). The system maintains provenance for both machine-generated and human-provided labels, allowing developers to track the source and evolution of document annotations (top dataframe in Figure 5). The developer may choose to display labels generated by the model or labels entered manually by an expert end-user, with the metadata system capturing the origin and timestamp of each label.
1@app.route("/")
2def home():
3 return flask.render_template("index.html")
4
5def get_colors():
6 infer = flor.dataframe("first_page", "page_color")
7 infer = flor.utils.latest(
8 infer[infer.document_value == pdf_names[-1]])
9 if infer.page_color.isna().any():
10 color = infer["first_page"].astype(int).cumsum()
11 infer["page_color"] = color - 1
12 return infer["page_color"].to_list()
13
14@app.route("/save_colors", methods=["POST"])
15def save_colors():
16 colors = request.get_json().get("colors", [])
17 pdf_name = pdf_names.pop()
18 with flor.iteration("document", None, pdf_name):
19 for i in flor.loop("page", range(len(colors))):
20 flor.log("page_color", colors[i])
21 flor.commit()
22 return jsonify({"message": "Colorssaved"}), 200
Figure 6. Flask routes demonstrating the human-in-the-loop feedback system: The home route serves the interface, get_colors() retrieves existing labels from FlorDB, and save_colors() captures expert corrections while maintaining provenance through FlorDB’s metadata management.
  • The root route (“/”) displays a home page which lists all the PDF files located in a specified directory. Each PDF file is represented with a preview image, and these images are listed on the webpage using rendered HTML templates.
  • The “/view-pdf” route handles requests to view a specific PDF. Depending on user interactions and the file’s existence, it can display the document in different modes such as labeled text or named entity recognition (NER) views. This route also supports expert annotation interfaces where users can correct model predictions.
  • The “/save_colors” route is a POST endpoint that processes user-submitted data concerning color settings associated with a PDF’s pages. This route captures this feedback data, logs it with appropriate metadata for tracking, and acknowledges the successful saving of data. The human feedback is stored with the same robust provenance tracking as computational results.
This bidirectional flow between computational processing and human expertise creates a complete feedback loop in the system. When experts provide corrections through the UI, the metadata system captures these annotations alongside the original model predictions. This human feedback can then be incorporated into subsequent model training iterations, with FlorDB maintaining clear provenance of which predictions were machine-generated versus human-corrected. This systematic approach to managing both computational and human feedback allows for continuous improvement of the model while maintaining transparency about the source and evolution of all labels in the system.
Takeaway: When used within human-in-the-loop interfaces, FlorDB functions as a comprehensive feedback management system (dataframe in Figure 5), maintaining provenance for both machine-generated and human-provided labels while enabling continuous model improvement through expert corrections.
@@ 5. Discussion
First, we assess how FlorDB embodies and extends the 3Vs of MLOps Shankar et al. (2024):
  • Velocity: FlorDB enhances the speed of ML development by getting metadata definition out of the critical path of agile experimentation. Rather than requiring upfront documentation that could slow iteration, FlorDB’s hindsight logging capability allows teams to move fast initially while preserving the ability to retroactively capture and analyze any needed metadata on demand. This approach maintains development velocity without sacrificing the rich context needed for effective ML lifecycle management.
@@
  • Versioning: FlorDB radically augments version control by providing a clean, simple interface to query metadata across versions, including technology for hindsight logging across those versions. This enables detailed tracking and allows teams to observe and understand arbitrary features of models as they evolve over time.
Building on these MLOps principles, FlorDB’s design draws inspiration from Ground’s vision for data context services Hellerstein et al. (2017). Like Ground, FlorDB recognizes that effective metadata management must capture not just static pre-declared descriptions but the full context of data usage across multiple dimensions. Specifically, FlorDB adapts Ground’s “ABCs of Data Context” - Applications, Behavior, and Change - but reframes them through an MLOps lens. While Ground focused broadly on data context for analytics, FlorDB specializes these concepts for the machine learning lifecycle, treating ML models and pipelines as first-class citizens that require rich contextual tracking.
5.1. Implications for Social Justice Research
FlorDB was developed in the context of the Berkeley EPIC Data Lab [^3] https://epic.berkeley.edu, whose mission encompasses democratizing data work via no-code and low-code interfaces, informed by applications in the social justice domain including criminal defense and investigative journalism. FlorDB’s design principles make it particularly well-suited for supporting social justice research efforts. By simplifying the tracking and organization of diverse datasets that social justice research often requires, FlorDB enables efficient data management even for users without extensive technical backgrounds. This accessibility is crucial for democratizing ML research capabilities across different communities and applications.
The system’s “metadata later” approach addresses key challenges faced by resource-constrained organizations, allowing teams to focus on urgent operational demands while maintaining the ability to refine documentation post-hoc as needed. For projects involving multiple data sources, such as public health studies or legal cases, FlorDB simplifies the tracking of data movement between sources, making processes more understandable to non-specialists Epstein and O’Brien (2019).
FlorDB’s focus on transparency and accountability (i.e. visibility) is especially valuable for investigating algorithmic bias and fairness. The system’s multiversion hindsight logging enables better tracking of decisions made during model training and development, helping researchers address concerns about the “black box” nature of ML models Rudin (2019). Moreover, its flexible metadata platform supports human-in-the-loop workflows, enabling continuous learning based on community feedback—a necessity for ethically deploying ML in dynamic social environments Holstein et al. (2019). By lowering technical barriers while enabling rigorous documentation and provenance tracking, FlorDB helps empower marginalized communities to participate more fully in ML research and development. This aligns with broader goals of democratizing access to ML technologies and ensuring their benefits extend equitably across society.
@@ 6. Related Work
Experiment Tracking and Version Control: Systems like MLFlow Zaharia et al. (2018), DVC dvc (2021), and Weights & Biases focus on managing experiments and ensuring reproducibility. They provide tools for tracking experiments, packaging code, and sharing models. While helpful for tracking model and data evolution, they primarily concentrate on experiment management without deeply integrating hindsight logging capabilities. As such, they do not (currently) address the challenges of context on demand that are the crux of our work.
@@
End-to-End ML Workflows: Systems like AWS SageMaker and Kubeflow provide comprehensive solutions for building, training, and deploying ML models Liberty et al. (2020); Community (2021). They offer tools for data labeling, model training, model hosting, and support for scalable ML workflows. Both platforms emphasize scalability and operational efficiency but primarily focus on deployment and operational aspects. FlorDB can use either system as a drop-in replacement to Make, the default build system. Other systems in this space include Helix Xin et al. (2018) and Motion Shankar and Parameswaran (2024).
@@
Data Catalogs and Metadata Management: Enterprise data catalogs like Collibra and Alation provide comprehensive solutions for organizing and discovering data assets across organizations. Collibra offers advanced governance features and automated workflows for policy management, while emphasizing data quality and observability capabilities Collibra, Inc. (2024). Alation focuses on intelligent data discovery through machine learning-powered search and active metadata management, helping organizations understand data lineage and trustworthiness Alation, Inc. (2024). In the open-source domain, DataHub and Amundsen represent modern approaches to metadata management. DataHub provides a platform spanning data discovery, observability, and governance LinkedIn, Inc. (2024), while Amundsen employs a modular architecture with dedicated services for metadata, search, and frontend interactions Lyft, Inc. (2024). These systems primarily focus on metadata discovery and governance, helping organizations create and maintain inventories of data assets across their digital landscape. While these platforms excel at data discovery and governance, they lack FlorDB’s deep integration with ML workflows and its ability to capture and recover fine-grained execution context.
@@ 7. Conclusion
FlorDB provides a new, general-purpose methodology and Python-based system for managing the complex metadata—the context—associated with the machine learning lifecycle. By resolving the longstanding friction between discipline and agility in metadata management, FlorDB offers AI/ML groups a flexible yet powerful solution aligning with modern MLOps agile development practices.
The key contributions of FlorDB include features that enhance the developer experience and streamline machine learning workflows. First, incremental context maintenance allows developers to gradually build out the metadata and project structure within their existing workflows. Second, the extensions presented to support pipelines, dataflow, and feedback are backward compatible with multiversion hindsight logging, allowing for a “metadata later” approach that enables the addition and refinement of metadata post-hoc. This capability supports rapid iteration and adaptation without compromising the long-term maintainability of projects. As demonstrated through the PDF Parser demo, FlorDB can take on multiple roles within the ML pipeline, acting as a feature store, model registry, training data store, and experiment record. This versatility eliminates the need for multiple specialized tools while maintaining robust metadata management throughout the project lifecycle.
The development of FlorDB was informed by an interview study of ML Engineers to understand how they operationalize the machine learning lifecycle Shankar et al. (2024), ensuring our design decisions were grounded in real-world practices and needs. While this evidence-based design approach helped shape FlorDB’s features and interface, a rigorous validation of its usability and other human factors through controlled user studies remains as important future work.
By providing a systematic way to balance agility and rigor, FlorDB creates new possibilities for scaling ML operations, improving collaboration between teams, and ensuring long-term maintainability of ML systems. As the field continues to evolve, FlorDB’s flexible architecture positions it to adapt to emerging needs while maintaining its core promise: enabling teams to move fast without sacrificing reproducibility and rigor.